# Improving Intrusion Detection with Domain-Invariant Representation Learning in Latent Space

### v2

@@ 3.2 The Multi-Task Learning Objective
In the unsupervised learning setting using latent space models, it is assumed that the label or the class information is not readily available. However, to achieve OOD generalization, we leverage the label information of the multiple source and cross-domain encoded data in the latent space during the training process.
In our approach, we do a joint optimization of the classification and the reconstruction loss using the latent space encoding. The total loss calculated over all the $N$ samples during the training process can be given as
$\begin{split}\mathcal{L}\left(\mathcal{S}_{train},Z;{\theta},{\phi}\right)=\min_{\phi,\theta}\frac{1}{N}\sum_{i=1}^{N}\Bigr{\{}\mathcal{L}^{\left(i\right)}_{ce}\left(g_{\phi}\left(X^{i}\right),y^{i}\right)\\ +\beta\mathcal{L}^{(i)}_{MI}\left(X^{i};Z^{i}_{\theta,\phi}\right)+\lambda\mathcal{L}_{rec}\left(X^{i};\theta,\phi\right)\Bigl{\}},\end{split}$ (6)
@@
$\mathcal{L}_{ce}^{(i)}\left(g_{\phi}\left(X^{i}\right),y^{i}\right)\\ =-\left(y^{i}\log\left(\mathcal{S}_{y}\left(g_{\phi}\left(X^{i}\right)\right)\right)\\ +\left(1-y^{i}\right)\log\left(1-\left(\mathcal{S}_{y}\left(g_{\phi}\left(X^{i}\right)\right)\right)\right)\right)$
$\mathcal{L}_{MI}$ is the disentanglement or decorrelation loss between the latent space $Z$ and the input space data $X$ in the form of mutual information as given in eq: 5, $\mathcal{S}_{y}$ is the softmax function applied on the encoded data $g_{\phi}(x)$, $\mathcal{L}_{rec}$ is the reconstruction loss. Here, we augment the cross-entropy loss with the mutual information regularization between the latent and prior space. The regularization aims to penalize spurious information content between the encoding and the prior space and the strength of the penalty is determined by the hyper-parameter $\beta$. $\beta$ controls the trade-off between decorrelation and classification. During the joint optimization, we try to balance the reconstruction weight and the mutual information regularization.
@@ 4.1 Dataset
  • CSE-CIC-IDS2018 This is a publicly available cybersecurity dataset that is made available by the Canadian Cybersecurity Institute (CIC). We combine the data from different days which has different classes of anomalies to get a multi-domain dataset. While testing, we consider the data from a different day which consists of other distinct classes of normal and anomalies as our OOD data.
  • NSL-KDD This is also a publicly available benchmark cybersecurity dataset made available by CIC. It has a total of 43 different features of internet traffic flow. The experimental results of our analysis are reported in our github repository.
@@ 4.2 Baselines
We consider the following models related to latent space multi-task representation learning and few-shot learning as baselines in order to compare and evaluate the performance of our proposed model.
  • Deep-CNN This is a supervised CNN model trained on all the features. The number of convolution, pooling, and fully connected layers are adjusted in order to achieve the best performance on all the classes. We perform an ensembling approach and calculate the average accuracy after running multiple iterations.
  • Deep Multi-task AE [1] In this latent dimension encoder-decoder-based model the reconstruction error is optimized over multiple domains in an unsupervised fashion. We consider multiple sources and cross-domain data and train them jointly with the label information in latent space. This is basically our model without the de-correlation regularization.
  • MMD-AE [3, 2] This paper uses the Minimum Mean Discrepancy (MMD) measure as regularization for domain generalization between multiple cross-domain data. We use it as a few-shot learning method where the cross-domain data are added to improve the OOD generalization.
  • Noise Enhanced Supervised AE (NSAE) [15]This model jointly predicts the labels of inputs and reconstructs the inputs considers reconstructed samples as noisy inputs and feeds them back to the model. In this model, it requires to do an extra step of fine-tuning by further training the reconstructed data using a supervised classifier.
  • MAML [11] For this baseline, we consider data samples from different domains to form the final classification task and evaluate the performance on unseen domain data (OOD). We consider the samples from the cross-domain classes to fine-tune the generalization process using the bi-level optimization algorithm.
@@ 4.3 Training Strategy
In the training phase, we jointly optimize multiple tasks by classifying multiple source and cross-domain samples from different datasets in the latent space. In our experiments with CIC-CSE-IDS data, our source domain data consists of the attack classes like GOLDEN-EYE, SOLARIS, and the BENIGN class, whereas our cross-domain samples are taken from attack classes like INFILTRATION, BRUTFORCE, and BOTNET. During the testing phase, we consider the OOD attack classes like DDOS-HOIC, RARE, and SlowHTTPTest along with an OOD BENIGN class to evaluate our domain generalization performance. During training, we introduce the cross-domain data in various proportions ranging from 5%-35% of the total source domain attack samples. This allows the model to learn a diverse latent space representation by reconstructing different source and cross-domain data while also simultaneously de-correlating spurious correlation between input and latent space by enforcing the mutual-information penalty. We use a learning rate of $0.005$ and train the model for $300$ epochs for all the training scenarios. We use a stacked encoder-decoder architecture which gradually projects the high-dimensional input space to low-dimensional latent space step-wise. The code is implemented in Pytorch and standard optimization libraries such as Adam are used during the training process.  [^1] https://anonymous.4open.science/r/MTRAE-40FA/
@@
     for batch = 1 to total batches do:
@@
         Compute the RBF kernels of prior space $K_{x_{l}}$ and latent space $K_{z_{l}}$
         Calculate mutual information between input space $X_{l}$ and latent space $Z_{l}$ using matrix-based Renyi’s entropy ${MI\left(X_{l}||Z_{l}\right)}$.
@@
         Calculate total batch loss $\mathcal{L}_{l}=\mathcal{L}_{ce}\left({X_{i},y_{i}}\right)+\mathcal{L}_{rec}\left({D\left(E\left(X_{i}\right)\right),X_{i}}\right)+\lambda\mathcal{L_{MI}}\left({X_{i}}||{Z_{i}}\right)$
@@
         $\mathbf{W}_{\phi_{t+1}}\leftarrow\mathbf{W}_{\phi_{t}}-\alpha_{1}\nabla_{\phi}\mathcal{L}\left(\theta,\phi\right)$
         $\mathbf{W}_{\theta_{t+1}}\leftarrow\mathbf{W}_{\theta_{t}}-\alpha_{2}\nabla_{\theta}\mathcal{L}\left(\theta,\phi\right)$
@@
Figure 1: Fig shows MI Gain($\beta$) variation with classification metrics on SOURCE and CROSS DOMAIN data (SOLARIS, GOLDENEYE, BENIGN,BRUT-FORCE, INFIL, BOTNET) and OOD DATA(RARE, HOIC, SLOW-HTTPS, BENIGN). We keep reconstruction loss weightage fixed. Fig shows that maximum accuracy is obtained when MI is between 1 and 3 for all the cases.
4.4 Sensitivity Analysis
In our joint optimization objective, we need to balance weightage to the hyperparameters $\beta$ and $\lambda$ in order to achieve good generalization performance to all the classes. We find that allocating a higher weightage to the entropy loss, $\beta=1.5$ to $\beta=3$, and keeping the reconstruction loss weightage at $\lambda=0.5$ to $\lambda=0.6$ yields the best performance of the model considering all the source, cross-domain, and OOD datasets. Fig 1 shows the maximum accuracy for most of the SOURCE, CROSS, and OOD domains achieved at this range. Another important hyper-parameter is the kernel bandwidth $\sigma$ used to estimate the matrix-based Renyi’s quadratic entropy in the latent space and prior space. We experiment with various kernel widths and report the best results in Table 1 and Table 2. We consider the batch size in between $200-300$ samples during each step of the gradient descent training as MI is easier to estimate when the sample size is relatively high.
ACCURACY-IN ACCURACY-OOD
@@ 4.4 Sensitivity Analysis
Model SOLARIS GOLD-EYE INFIL BOTNET RARE HOIC SLOW-HTTPS BENIGN
MAML $93.50$ $99.70$ $69.14$ $73.50$ $30.31$ $86.03$ $54.24$ $34.58$
Deep-CNN $\mathbf{100.00}$ ${\mathbf{100.00}}$ 99.0599.05{99.05}99.05 50.9550.95{50.95}50.95 $36.90$ $89.16$ $58.56$ $71.34$
D-MTAE $99.50$ $99.60$ $\mathbf{99.40}$ $99.50$ $\mathbf{61.30}$ 27.3827.38{27.38}27.38 $32.20$ $61.00$
MMD-AAE $99.34$ $99.70$ 99.0499.04{99.04}99.04 $99.11$ $48.05$ $28.00$ 02.0002.00{02.00}02.00 $46.25$
NSAE $\mathbf{100.00}$ $\mathbf{100.00}$ $55.07$ 50.0050.00{50.00}50.00 48.7648.76{48.76}48.76 $\mathbf{100.00}$ $50.56$ 61.3961.39{61.39}61.39
MTLS-RED 88.8388.83{88.83}88.83 86.0886.08{86.08}86.08 82.4182.41{82.41}82.41 $\mathbf{96.12}$ 28.0928.09{28.09}28.09 80.9980.99{80.99}80.99 $\mathbf{73.93}$ $\mathbf{73.23}$
Table 1: Table 1 shows the accuracy of different SOURCE domains comprising of attack classes -BENIGN, Solaris, and Golden-Eye, the CROSS DOMAIN attacks Infiltration, SSH, and BOTNET. The OOD classes are RARE, SlowHTTPTest, HOIC, and OOD BENIGN. In this case, we add 10% of source domain attack samples as cross-domain data.
ACCURACY-IN ACCURACY-OOD
Model SOLARIS GOLD-EYE INFIL BOTNET RARE HOIC SLOW-HTTPS BENIGN
MAML $97.35$ $97.40$ $54.84$ $87.51$ $46.39$ $70.45$ $34.78$ $70.45$
Deep-CNN $\mathbf{100.00}$ $\mathbf{100.00}$ $\mathbf{99.94}$ 51.4151.41{51.41}51.41 $36.92$ $61.65$ $49.52$ $77.81$
D-MTAE $61.30$ $61.30$ $32.20$ $61.00$ $72.60$ $48.40$ $69.80$ $\mathbf{83.32}$
MMD-AAE $96.61$ $99.46$ 99.2199.21{99.21}99.21 $78.23$ $24.84$ $72.08$ 47.1347.13{47.13}47.13 $46.25$
NSAE 99.9899.98{99.98}99.98 $\mathbf{100.00}$ $34.85$ 49.4249.42{49.42}49.42 48.4048.40{48.40}48.40 $\mathbf{82.70}$ $\mathbf{99.25}$ 51.5951.59{51.59}51.59
MTLS-RED $79.46$ $79.33$ $79.20$ $\mathbf{79.12}$ $\mathbf{78.99}$ 79.3379.33{79.33}79.33 $79.25$ 79.6479.64{79.64}79.64
Table 2: Table 2 shows the accuracy of the SOURCE data comprising of attack classes -BENIGN, Solaris, and Golden-Eye, the CROSS DOMAIN attacks Infiltration, SSH, and BOTNET. The OOD classes are RARE, SlowHTTPTest, HOIC, and OOD BENIGN. In this case, we add 35% of source domain attack samples as cross-domain data.
@@ 4.5 Ablation Study
The ablation study is performed keeping the baseline models as reference. In order to study the effectiveness of our proposed model, we perform the following analysis
  • We consider the multiple source and cross-domain reconstructions and use the encoded space to classify SOURCE, CROSS-DOMAIN, and OOD data. Our setup is different from [1] where the authors adopt a two-stage approach where the model is trained first, and then the trained encoder is used for classification purposes. On the other hand, our strategy is to jointly train all the components together so that the joint optimization with the de-correlation penalty can improve the generalization performance. The joint objective promotes discriminative power while simultaneously reducing the dependency on spurious dataset-specific correlation.
  • In our experiments we increase the cross-domain data availability gradually from 5%-35% in the training data in order to understand the impact of decorrelation penalty in achieving domain generalization while reconstructing data from multiple domains. In Table 1, we observe that when the cross-domain attack classes constitute a small percentage $10\%$ of the source domain data, most of the baseline models overfit the source domain classes and do not perform well on the cross-domain and test OOD classes. The OOD classes are never involved in the training process. Whereas, our proposed model gives an overall good performance on all the source, cross-domain, and OOD classes except the RARE class. In Table 2, we report the results of the experiments where we gradually increase the cross-domain data to $35\%$. In this case, with our proposed model, the best accuracy generalizes well to all the source, cross-domain, and OOD classes although there is a decrease in performance with some of the source and cross-domain classes but better performance can be seen with the OOD classes especially. In a nutshell, we want to point out that by adjusting the kernel bandwidth and the sample size, we can get a reliable estimate of the mutual information penalty between prior and latent space which can help achieve better generalization to completely unseen classes that were never used in the training process.

### v3

@@ 3.2 The Multi-Task Learning Objective
In our latent space multi-task learning approach, we leverage the label information of the multiple source and cross-domain encoded data in the latent space during the training process. In our approach, we do a joint optimization of the classification and the reconstruction loss along with the MI regularization in the latent space. The total loss calculated over all the $N$ batches, each consisting of $l$ samples, during the training process can be given as
$\begin{split}\mathcal{L}\left(\mathcal{S}_{train},Z;{\phi},{\theta}\right)=\min_{\phi,\theta}\frac{1}{N}\sum_{l=1}^{N}\Bigr{\{}\mathcal{L}_{ce}\left(g_{\phi}\left(X^{l}\right),y^{l}\right)\\ +\beta\mathcal{L}_{MI}\left(X^{l};Z^{l}\right)+\lambda\mathcal{L}_{rec}\left(X^{l};\phi,\theta\right)\Bigl{\}},\end{split}$ (6)
@@
$\mathcal{L}_{ce}\left(g_{\phi}\left(X\right),y\right)\\ =-\left(y\log\left(\mathcal{S}_{y}\left(g_{\phi}\left(X\right)\right)\right)\\ +\left(1-y\right)\log\left(1-\left(\mathcal{S}_{y}\left(g_{\phi}\left(X\right)\right)\right)\right)\right)$
$\mathcal{L}_{MI}$ is the disentanglement or de-correlation loss between the latent space and the input space expressed in the form of mutual information divergence measured in their kernel space, given in eq: 5, $\mathcal{S}_{y}$ is the softmax function applied on the encoded data $g_{\phi}(x)$, $\mathcal{L}_{rec}$ is the reconstruction loss, $\phi$, $\theta$ are the encoder and decoder parameters. Here, we augment the cross-entropy loss with the mutual information regularization between the latent and prior space. The regularization aims to penalize spurious information content between the latent encoding and the input space, the strength of which is determined by the hyper-parameter $\beta$. $\beta$ controls the trade-off between decorrelation and classification. During the joint optimization, we try to balance the reconstruction weight and the mutual information regularization.
@@ 4.1 Dataset
  • CSE-CIC-IDS2018 This is a publicly available cybersecurity dataset that is made available by the Canadian Cybersecurity Institute (CIC). It consists of 7 major kinds of intrusion datasets. We use SOLARIS, GOLDENEYE as source domain data, INFILTRATION, BOTNET as cross-domain data, and RARE, SLOWHTTPS, HOIC as OOD data, along with a OOD BENIGN class from a different day.
  • CICIoT 2023 This is a state-of-the-art dataset for profiling, behavioral analysis, and vulnerability testing of different IoT devices with different protocols from the network traffic, consisting of 7 major attack classes. We use BENIGN, DoS, and DDoS as source data, RECON, as cross-domain data and WEB, MIRAI as OOD test data.
  • CICIoMT 2024 This is a benchmark dataset to enable the development and evaluation of Internet of Medical Things (IoMT) security solutions. The attacks are categorized into five classes. We use BENIGN, DDoS, DoS as source-domain, RECON, and SPOOFING as cross-domain, and MQTT as OOD data.
 [^1] CIC- Canadian Institute of Cybersecurity
@@ 4.2 Baselines
We consider the following models related to latent space multi-task representation learning and few-shot learning as baselines to compare and evaluate the performance of our proposed model. More details about the baselines can be found in the supplementary material.
  • Correlation Alignment for Deep Domain Adaptation (CORAL) [39] This paper adopts an unsupervised domain adaptation technique that tries to align the covariance of the source and the target distributions. We extend CORAL to latent space and align the source and cross-domain data correlation to improve generalization on OOD data.
  • Deep Multi-task Autoencoder (MTAE) [1] In this latent dimension encoder-decoder-based model the reconstruction error is optimized over multiple domains in a supervised fashion. We consider multiple sources and cross-domain data and train them jointly with the label information in latent space. This is basically our model without the de-correlation regularization.
  • Minimum Mean Discrepancy-Autoencoder(MMD-AE)[3, 2] This paper uses the MMD measure as regularization for domain generalization between multiple cross-domain data. We use it as a few-shot learning method where the cross-domain data are added to improve the OOD generalization.
  • Noise Enhanced Supervised Autoencoder (NSAE) [15]This model jointly predicts the labels of inputs and reconstructs the inputs considers reconstructed samples as noisy inputs and feeds them back to the model. In this model, requires doing an extra step of fine-tuning by further training the reconstructed data using a supervised classifier.
@@ 4.3 Training Strategy
In the training phase, we jointly optimize multiple tasks by considering multiple source and cross-domain data in the latent space. We consider a setting where the joint distribution of the data doesn’t have much overlap but the marginal distribution of the features do have an overlap. This scenario happens when the features exhibit unique correlation patterns between the source and cross-domain data distributions. For example, DOS and DDOS datasets have similar correlation patterns between the features. However, the the features in the other attacks such as INFIL and WEB have different correlation patterns. Similarly, the GOLDEN and SOLARIS datasets have a certain kind of correlation between the features that differ from other attack datasets like INFIL and BOTNET. Our goal is to combine the datasets with different patterns of correlation structure among its features during the training phase in order to increase the versatility of the latent space. For example, if two or more features of any two attack datasets exhibit distinct correlation patterns, we add one as the source domain and the other as cross-domain data. This will tend to introduce more variety in the input space and more generalized information in the latent space after applying the mutual information penalty. During training, we vary the cross-domain data in various proportions ranging from 5%-50% of the source domain samples and evaluate the results on the OOD data. This allows the model to learn a diverse latent space representation by reconstructing different source and cross-domain data while simultaneously de-correlating domain-specific spurious correlation information.
Figure 1: Fig (a-e) Correlation plots of various cyber-security datasets. (a)DOS (b) DDOS (c) WEB (d) INFIL (e) HOIC
 [^2] https://anonymous.4open.science/r/MTRAE-40FA/
@@
     for batch = 1 to total batches N do:
@@
         Compute the RBF kernels of input space $\mathcal{K}_{x_{l}}$ and latent space $\mathcal{K}_{z_{l}}$ of size $l\times l$
         Calculate mutual information between input space $X_{l}$ and latent space $Z_{l}$ using matrix-based Renyi’s entropy ${MI\left(\mathcal{K}_{{X}_{l}};\mathcal{K}_{{Z}_{l}}\right)}$.
@@
         Calculate total batch loss $\mathcal{L}_{l}=\mathcal{L}_{ce}\left({X^{l},y^{l}}\right)+\mathcal{L}_{rec}\left(X^{l},X^{{l}^{\prime}}\right)+\lambda\mathcal{L_{MI}}\left({X}^{l}||{{Z}^{l}}\right)$
@@
         $\mathbf{W}_{\phi_{t+1}}\leftarrow\mathbf{W}_{\phi_{t}}-\alpha_{1}\nabla_{\phi}\mathcal{L}_{l}\left(\theta,\phi\right)$
         $\mathbf{W}_{\theta_{t+1}}\leftarrow\mathbf{W}_{\theta_{t}}-\alpha_{2}\nabla_{\theta}\mathcal{L}_{l}\left(\theta,\phi\right)$
@@
4.4 Hyperparameter Sensitivity
In the joint optimization objective, we aim to balance the weightage to the hyperparameters $\beta$ and $\lambda$ to achieve good generalization performance to all the classes. We find that allocating a higher weightage to the entropy loss and keeping less weightage to the reconstruction loss results in the best overall performance of the model considering different scenarios. The kernel bandwidth used to estimate the matrix-based Renyi’s quadratic entropy in the latent space and input space proves to be a significant hyperparameter. Optimum bandwidth selection is directly linked to the smoothness of the learning. A small bandwidth can lead to under-smoothing whereas a large bandwidth may result in over-smoothing We vary the kernel bandwidth during our experiments. We consider the correlated features in the case of each dataset. For a fair comparison between the baselines, we keep the latent space dimension similar while experimenting on the datasets with a particular baseline.
Datasets Features learning rate kernel bw batch size Weightage Latent Dim
CSE-CIC-IDS 79 0.005 0.01-6 200 $\lambda,\beta=0.6,2$ $79-30-15$
CICIoT 2022 43 0.005 0.01-6 200 $\lambda,\beta=0.6,2$ $43-27-13$
CICIoMT 2024 44 0.005 0.01-6 200 $\lambda,\beta=0.6,2$ $44-22-11$
SOURCE DOMAIN CROSS DOMAIN OOD DOMAIN
@@ 4.4 Hyperparameter Sensitivity
Model Percentage SOLARIS GOLD-EYE INFIL BOTNET RARE HOIC HTTPS BENIGN
DMTAE $0\%$ $\mathbf{99.97}$ 52.7052.70{52.70}52.70 04.0004.00{04.00}04.00 09.5009.50{09.50}09.50 $00.30$ $00.38$ $02.30$ 97.9897.98{97.98}97.98
$15\%$ 99.5099.50{99.50}99.50 $\mathbf{99.70}$ $\mathbf{99.40}$ $\mathbf{99.50}$ $61.30$ $27.38$ $32.30$ $\mathbf{98.11}$
$25\%$ 81.6081.60{81.60}81.60 79.6079.60{79.60}79.60 69.7069.70{69.70}69.70 78.5078.50{78.50}78.50 $65.40$ $\mathbf{50.00}$ $42.40$ $62.00$
$40\%$ 61.6161.61{61.61}61.61 $61.30$ 31.3031.30{31.30}31.30 $61.60$ $\mathbf{72.60}$ $48.50$ $\mathbf{69.80}$ 83.3283.32{83.32}83.32
MMD-AE $0\%$ $\mathbf{99.99}$ $\mathbf{99.29}$ 00.5300.53{00.53}00.53 $\mathbf{99.98}$ $\mathbf{55.47}$ $\mathbf{99.98}$ $\mathbf{99.69}$ $\mathbf{99.71}$
$15\%$ 99.9899.98{99.98}99.98 02.5502.55{02.55}02.55 $\mathbf{22.55}$ 99.8399.83{99.83}99.83 12.1912.19{12.19}12.19 $41.34$ $98.77$ $98.67$
$25\%$ $99.78$ 02.5202.52{02.52}02.52 $15.35$ $99.83$ $12.19$ $41.27$ $99.69$ $98.61$
$40\%$ $99.96$ 01.8801.88{01.88}01.88 12.3012.30{12.30}12.30 $99.83$ $14.13$ 41.1241.12{41.12}41.12 56.2156.21{56.21}56.21 99.1499.14{99.14}99.14
NSAE $0\%$ 99.9999.99{99.99}99.99 99.9899.98{99.98}99.98 00.0900.09{00.09}00.09 99.9999.99{99.99}99.99 $\mathbf{77.03}$ $\mathbf{99.90}$ $\mathbf{97.02}$ 99.5899.58{99.58}99.58
$15\%$ 99.8099.80{99.80}99.80 $\mathbf{99.99}$ $04.82$ $34.92$ $36.04$ $00.00$ $05.16$ $\mathbf{99.80}$
$25\%$ $\mathbf{99.99}$ $03.16$ 00.3200.32{00.32}00.32 $\mathbf{99.83}$ $59.36$ $15.90$ $00.35$ $99.69$
$40\%$ 99.9899.98{99.98}99.98 34.3634.36{34.36}34.36 $\mathbf{53.66}$ $99.83$ 12.3612.36{12.36}12.36 00.0000.00{00.00}00.00 19.3319.33{19.33}19.33 $99.40$
CORAL $0\%$ 59.1859.18{59.18}59.18 90.6190.61{90.61}90.61 30.4530.45{30.45}30.45 00.8000.80{00.80}00.80 $08.40$ $55.06$ $03.40$ $69.94$
$15\%$ $61.53$ $12.64$ $\mathbf{31.79}$ $50.43$ $\mathbf{33.74}$ $41.31$ 50.3850.38{50.38}50.38 $67.54$
$25\%$ $38.85$ $\mathbf{99.99}$ $0.00$ $0.01$ 22.9622.96{22.96}22.96 $82.21$ $0.00$ $95.01$
$40\%$ $\mathbf{99.99}$ $38.85$ 00.0100.01{00.01}00.01 00.0000.00{00.00}00.00 $22.96$ $\mathbf{82.21}$ 00.0100.01{00.01}00.01 $\mathbf{99.19}$
MTLS-RED $0\%$ $\mathbf{88.83}$ $\mathbf{86.08}$ $\mathbf{82.41}$ $\mathbf{96.12}$ $28.01$ $80.99$ $73.93$ $73.23$
$15\%$ 79.0079.00{79.00}79.00 78.8578.85{78.85}78.85 78.7178.71{78.71}78.71 78.6878.68{78.68}78.68 $78.88$ $79.13$ $78.82$ $79.01$
$25\%$ $83.90$ $80.70$ $80.70$ $83.80$ $76.50$ $\mathbf{81.96}$ $\mathbf{85.00}$ $77.10$
$40\%$ 79.4679.46{79.46}79.46 79.3379.33{79.33}79.33 79.2079.20{79.20}79.20 79.1279.12{79.12}79.12 $\mathbf{78.99}$ 79.3379.33{79.33}79.33 79.2579.25{79.25}79.25 $\mathbf{79.64}$
Table 1: Performance of the proposed method and baseline method with different percentages of cross-domain data added to the source domain during training. The results are on the CIC-CSE-IDS dataset. We highlight the best performance of each model across all the datasets.
@@ 4.5 Ablation Study
We perform different kinds of ablation studies to validate the performance of our model and also to compare our proposed model with standard baselines. In our experiments, we consider multiple source and cross-domain reconstructions and use the encoded space to classify source, cross-domain, and OOD data. Our setup is different from MTAE [1] where the authors adopt a two-stage approach where the model is trained first, and then the trained encoder is used for classification purposes. On the other hand, our strategy is to jointly optimize all the components together to improve the generalization performance. The joint objective promotes discriminative power while simultaneously reducing the dependency on spurious data-specific correlation. We try to design a principled approach to how cross-domain data can be combined with particular source domain data in order to improve the generalization to out-of-distribution data samples. We vary the cross-domain data in various proportions 5%-50% of the source data in the training data in order to understand the impact of the de-correlation penalty under different scenarios. In Table 1,2, we observe that as we gradually increase the percentage of the cross-domain classes in the training data, the performance of most of the baseline models deteriorates on the IN distribution. Whereas, with a small percentage of cross-domain data initially, most of the models overfit to the IN distribution datasets. The proposed model, on the other hand, gives an overall good performance on all the source, cross-domain, and OOD classes as we vary the amount of cross-domain data. It is to be noted that the OOD-domain data are never observed during any stage of the training process. In Table 1, we observe an interesting phenomenon where we gradually increase the cross-domain data to $30\%-40\%$, by adjusting the kernel bandwidth and the hyperparameters $\beta$ and $\lambda$, we can train a model that shows an overall generalization performance across all training and test datasets. The base models mostly overfit to the training data or a particular cross-domain data. In Table 2, we find that when we increase the proportion of cross-domain data during the training, the classification results are significantly good across the OOD datasets like WEB and SPOOFING attacks. Although there is a decrease in performance in some of the source domains, an overall improvement in performance across all classes can be seen. Determining the value of the kernel bandwidth parameter in kernel density estimation is important as it leads to over-smoothing or under-smoothing. In Figure 2, we plot the latent space of a simple multi-task encoder-decoder model without any regularization and our proposed model with the MI penalty. We see improved clustering with many of the source, cross, and target domain classes when the regularization is applied.
Figure 2: Fig (f-j),(p-t) T-SNE projection of the latent space of DMTAE (without MI regularization) and (a-e),(k-o) is our proposed MTL-RED (with MI regularization). Again, (a-j) the attacks are GOLDEN-EYE, SOLARIS, BOTNET, HOIC, RARE. (k-t) the attacks are DOS, DDOS, MQTT, RECON, and SPOOFING. We see the difference distinctively in SOLARIS(b,g), RARE(e,j), DOS(k,p), DDOS(l,q) and MQTT(m,r).
SOURCE DOMAIN CROSS DOMAIN OOD DOMAIN
Model Percentage DDOS DOS RECON SPOOF MQTT MIRAI WEB BENIGN
DMTAE $0\%$ $99.96$ 99.9999.99{99.99}99.99 54.0754.07{54.07}54.07 46.9846.98{46.98}46.98 $78.43$ $\mathbf{99.97}$ $30.55$ $97.53$
$15\%$ $99.99$ 99.9999.99{99.99}99.99 98.6698.66{98.66}98.66 71.8471.84{71.84}71.84 $78.43$ $79.38$ $\mathbf{30.55}$ $97.53$
$25\%$ 99.9999.99{99.99}99.99 99.9999.99{99.99}99.99 $\mathbf{99.99}$ $\mathbf{75.55}$ $\mathbf{99.93}$ $80.07$ $21.38$ $98.63$
$40\%$ $\mathbf{99.99}$ $\mathbf{99.99}$ 98.53498.534{98.534}98.534 $74.43$ $89.42$ $80.83$ $29.02$ $97.78$
MMD-AE $0\%$ $99.96$ $99.96$ $49.95$ $41.32$ $\mathbf{95.87}$ $84.01$ $21.15$ $\mathbf{98.56}$
$15\%$ $\mathbf{99.99}$ $\mathbf{99.99}$ $98.47$ $70.54$ $90.54$ $\mathbf{84.01}$ $42.53$ $96.49$
$25\%$ $99.99$ $99.99$ $99.11$ $73.60$ $89.42$ $77.60$ $58.04$ $93.48$
$40\%$ $99.61$ $99.46$ $\mathbf{99.30}$ $\mathbf{78.23}$ $94.84$ $72.08$ $\mathbf{67.21}$ $92.41$
NSAE $0\%$ $\mathbf{99.99}$ $99.99$ $46.90$ $32.42$ $69.11$ $\mathbf{99.90}$ $\mathbf{56.05}$ $94.68$
$15\%$ $99.99$ $\mathbf{99.99}$ $97.77$ $69.11$ $\mathbf{99.90}$ $99.87$ $36.66$ $\mathbf{97.71}$
$25\%$ $99.99$ $99.99$ $\mathbf{98.70}$ $\mathbf{71.29}$ $99.71$ $99.90$ $43.83$ $96.11$
$40\%$ $99.99$ $99.99$ $98.62$ $70.64$ $99.89$ $99.90$ $49.70$ $95.78$
CORAL $0\%$ $\mathbf{99.99}$ $99.99$ $98.53$ $74.43$ $89.42$ $\mathbf{81.13}$ $30.53$ $98.73$
$15\%$ $99.99$ $\mathbf{99.99}$ $98.53$ $74.43$ $89.42$ $81.13$ $\mathbf{42.30}$ $98.73$
$25\%$ $99.99$ $99.99$ $\mathbf{98.86}$ $75.55$ $\mathbf{99.90}$ $80.08$ $42.03$ $\mathbf{98.73}$
$40\%$ $99.99$ $99.99$ $99.11$ $\mathbf{77.21}$ $99.57$ $80.51$ $42.30$ $98.73$
MTLS-RED $0\%$ $99.99$ $99.99$ $98.41$ 78.9278.92{78.92}78.92 $78.21$ 76.0476.04{76.04}76.04 $73.91$ $87.67$
$15\%$ $99.99$ $99.99$ $98.41$ 78.9278.92{78.92}78.92 $78.21$ 99.7599.75{99.75}99.75 $68.03$ $91.83$
$25\%$ $99.99$ $99.99$ $98.41$ $78.21$ $53.19$ 99.8799.87{99.87}99.87 $73.67$ $93.05$
$40\%$ $\mathbf{99.99}$ $\mathbf{99.99}$ $\mathbf{98.98}$ $\mathbf{81.17}$ $\mathbf{95.87}$ $\mathbf{99.88}$ $\mathbf{75.91}$ $\mathbf{91.2}$
Table 2: Performance of the proposed method and baseline method with different percentages of cross-domain data added to the source domain during training. The results are on the combined features of CIC-IOT and CIC-IoMT datasets.